# Federated Unsupervised Domain Generalization using Global and Local Alignment of Gradients

### v1

@@ Abstract
We address the problem of federated domain generalization in an unsupervised setting for the first time. We first theoretically establish a connection between domain shift and alignment of gradients in unsupervised federated learning and show that aligning the gradients at both client and server levels can facilitate the generalization of the model to new (target) domains. Building on this insight, we propose a novel method named FedGaLA, which performs gradient alignment at the client level to encourage clients to learn domain-invariant features, as well as global gradient alignment at the server to obtain a more generalized aggregated model. To empirically evaluate our method, we perform various experiments on four commonly used multi-domain datasets, PACS, OfficeHome, DomainNet, and TerraInc. The results demonstrate the effectiveness of our method which outperforms comparable baselines. Ablation and sensitivity studies demonstrate the impact of different components and parameters in our approach. The source code will be available online upon publication.
@@ 1 Introduction
Federated learning [1, 2] has emerged as a promising framework for training machine learning models across multiple decentralized clients while preserving data privacy. It allows clients to collaboratively train a global model without the need to exchange their sensitive and local data. Each client trains a local model using its data and a server aggregates these models at a certain frequency [3, 4, 5]. However, given that each client collects a different set of local training data, two issues arise. First, the data collected by each client is often recorded under unique conditions that may result in mutual domain shifts [6]. Second, labeling training data is inherently challenging and resource-intensive; this issue is even more pronounced in the context of federated settings where different clients may resort to different labeling strategies and protocols [7].
In prior works, each of these issues has been addressed as a separate problem statement: (i) federated domain generalization [8, 9, 10], and (ii) federated unsupervised learning [7, 11]. Despite the effectiveness of both problem definitions, each ignores the fundamental assumptions of the other regarding the data in terms of distributions and availability of labels. To further approach federated learning in a more practical and real-world scenario, we propose to merge these two under a new umbrella called federated unsupervised domain generalization, which we define as Definition 1. To our knowledge, prior works have not studied federated learning under such constraints.
@@
To address this new problem, we propose a novel method called Federated Unsupervised Domain Generalization using Global and Local Alignment of Gradients (FedGaLA). To learn more generalized representations from multiple domains, FedGaLA relies on gradient alignment at both client (local) and server (global) levels. At the client level, local models are trained with unlabeled local data available to each client using self-supervised learning (SSL). To learn domain-invariant representations, gradients that are not aligned with the reference gradient, i.e., the global learning direction, are discarded. At the server level, to achieve better generalization, the local models are aggregated based on their alignment with each other (see Figure 1).
@@
Specifically, the local gradients that are more closely aligned with the average gradient are given greater weight in the aggregation stage. Moreover, we provide a detailed theoretical framework establishing a connection between the alignment of gradients from different clients and the similarity between their data distributions, and thus better domain generalization. Since FedGaLA does not require the sharing of data between clients to perform alignment at either the server or client sides, it effectively preserves data privacy. We verify the performance of our approach using four public datasets, PACS [12], OfficeHome [13], DomainNet [14], and TerraInc [15] and demonstrate strong performance in federated unsupervised domain generalization in comparison to various baselines. Additionally, we conduct different ablation and sensitivity studies to understand the impact of different parameters and the choice of SSL frameworks.
In summary, the main contributions of this study are as follows. (1) To our knowledge, this is the first work to address federated unsupervised domain generalization, introducing a new problem class in this area. (2) We propose a novel technique that considers and aligns gradients at both the local and global levels. Doing so, our solution effectively extracts domain-invariant information in local training and aligns each client’s contributions during the aggregation process, boosting the model’s generalization across various target domains. (3) We conduct extensive experiments and ablations to demonstrate the effectiveness of our proposed method on various benchmarks. We will make our code public upon publication.
@@ 2 Related work
Federated learning. Federating learning is a technique of distributed training that enables learning from decentralized clients without the need to transfer raw data to a central server [1, 2]. In the first work in the area of federated learning, FedAVG [1] aggregates the weights of clients via averaging to form the global model. Various federated learning techniques have been since developed for different purposes. FedProx [16] adds a proximal term to the objective to tackle the data heterogeneity problem. FedBuf [17] proposes to use buffered asynchronous aggregation to address scalability and privacy issues. FedALA [18] dynamically aggregates the downloaded global model with the local model on each client, ensuring alignment with the local objective. Moon [5] leverages the similarity between model representations to improve the local training of individual parties through client-level contrastive learning.
@@
Federated domain generalization. This category of problems aims to learn a general model in a privacy-preserving manner that generalizes well to new target domains with distribution shifts [6, 19]. There have been several studies to explore this direction, mainly focusing on learning domain-invariant features [9, 20] or identifying common features across multiple domains [8]. FedDG [6] allows clients to share their data in the frequency space for medical image segmentation. FedSR [9] proposes to learn a simple representation of data to better generalize to target domains through two regularizers, i.e., L2-norm and conditional Mutual Information. FedADG [8] leverages adversarial training to align source domain distributions by matching each distribution with a reference distribution. FedKA [21] learns domain-invariant features by employing feature distribution matching in a universal workspace. CCST [22] aligns various client distributions and mitigates model biases by adapting local models to diverse sample styles via cross-client style transfer. StableFDG [23] uses style and attention-based strategies to address the federated domain generalization problem. Finally, FedDG-GA [24] uses domain divergence and a moment mechanism to enhance generalization through dynamic domain weight adjustment.
@@
Federated unsupervised learning. This category attempts to learn representations from unlabeled data distributed across clients while preserving data privacy [25]. The study by [26] is the first to propose federated unsupervised learning using an encoder-decoder structure. FedU [7] employs a contrastive learning approach with online and target networks to enable each client to independently learn representations from unlabeled data. It also introduces a dynamic aggregation mechanism to update the predictor, either locally or globally. Similarly, FedEMA [27] uses the exponential moving average of the global model to update local ones. FedX [11] proposes to use knowledge distillation to learn representation from local data and refine the central server’s knowledge. FedCA [28] consists of two components: (i) a dictionary module to gather the representations of samples from each client to maintain consistency in the representation space across all clients, and (ii) an alignment module to adjust each client’s representation to match a base model trained on public data.
Federated unsupervised domain generalization. As discussed earlier in Section 1, we define a new problem space not studied previously in the literature, which unifies federated domain generalization and federated unsupervised learning. The key distinction between our new problem definition and standard federated domain generalization is that we assume labels are not available at the client or server levels. On the other hand, our new problem definition differs from standard federated unsupervised learning in that our problem definition assumes domain shifts among clients while existing federated unsupervised methods assume identical data distributions among all clients.
@@ 3.1 Problem formulation
To formalize Definition 1, assume $K$ clients, $C_{i}$, in a federated setup, each with access to its own unlabeled data $\mathcal{D}_{i}=\{\textbf{x}_{i}^{(n)}\}_{n=1}^{N_{i}}$. Each dataset consists of $N_{i}$ data points sampled from a distinct data distribution $p(\textbf{x}_{i})$, where $\textbf{x}_{i}$ is a vector of $F$ features, i.e., $\textbf{x}_{i}=[x_{i}^{1},x_{i}^{2},...x_{i}^{F}]^{T}$. The data distributions are assumed to be different among the clients with each distribution $p(\textbf{x}_{i})$ sampled from a family of distributions $\mathcal{P}$. Privacy constraints prevent the transfer of data between clients or to the server $S$. The objective is to learn generalized representations from $\textbf{x}_{i}$ that perform well across unseen distributions $p(\textbf{x}_{t})\sim\mathcal{P}$, where $p(\textbf{x}_{i})\neq p(\textbf{x}_{t})$. This is formulated as minimizing the expected loss over the unseen distributions:
@@
$\min_{\theta_{i}}\mathbb{E}_{p(\textbf{x}_{i})}\left[l_{i}(\theta_{i};\textbf{x})\right]\approx\frac{1}{N_{i}}\sum^{N_{i}}_{n=1}l(\theta_{i};\textbf{x}_{i}^{(n)}),$ (2)
@@ 3.2 Gradient alignment and domain shift
Under a federated learning framework, privacy constraints prevent clients and servers from accessing each other’s data, including distribution information such as data means and variances. They can, however, observe individual client gradients at the server level and the average aggregated gradient across clients at the client level. We motivate our work on the fact that alignment of gradients may infer characteristics of the client domain distributions, thus facilitating improved model generalization. While empirically the utility of gradients has been demonstrated in the area of domain generalization [29], no theoretical basis has been proposed for this approach under federated constraints. Accordingly, we aim to establish a link between gradient alignment and domain shifts within the proposed problem formulation (Definition 1). Our theoretical findings provide the basis for effective local parameter updates and global model aggregation to address federated unsupervised domain generalization.
@@
Let each $\textbf{x}_{i}^{f}$ be a random variable drawn from a Normal distribution $p(\textbf{x}_{i}^{f})\sim\mathcal{P}$ [30]. Within a single domain, following [31, 32], features are assumed to be independent ($Cov(x_{i}^{f_{1}},x_{i}^{f_{2}})=0$). Across different domains, corresponding features of $\textbf{x}_{i}^{f}$ and $\textbf{x}_{j}^{f}$ are bivariate with a covariance of $\sigma_{x_{i}^{f},x_{j}^{f}}$. In line with contemporary practices in deep learning, the features are normalized with $\mu=0$ and $\sigma^{2}=1$ [33, 34, 35]. We also assume that each client is a one-layer encoder with the sigmoid activation function. Following [36], local training is performed using contrastive loss where a random augmentation of a data sample is used as the positive pair and all other samples are utilized as the negative pairs. The random augmentation is performed with the same random Affine transformation ($Ax_{i}+B$) [37] broadcast to all clients. After each epoch, the local models are aggregated in the server and sent back to all clients. The gradient $\textbf{g}_{i}$ of the model is assumed to be differentiable.
@@
Given Assumption 1, under the problem proposed in Definition 1, for two distinct domains characterized by random variables $\textbf{x}_{i}$ and $\textbf{x}_{j}$ belonging to two different clients $C_{i}$ and $C_{j}$, an increase in domain shift across the clients results in a decrease in the covariance $\text{Cov}(\emph{$\textbf{g}_{i}$},\emph{$\textbf{g}_{j}$})$ of the corresponding gradients $\textbf{g}_{i}$, $\textbf{g}_{j}$ across $C_{i}$ and $C_{j}$’s respective local models.
@@
Motivated by previous works where the similarity between the representations of different domains under domain shift is measured by Mutual Information [38, 39, 40], we use this concept for modeling the similarity between different domains drawn from a family of distributions. By definition, the Mutual Information between two random variables, $\textbf{x}_{i}$ and $\textbf{x}_{j}$ is
[…]
$p(\textbf{x}_{i},\textbf{x}_{j})=\prod_{f=1}^{F}p(x_{i}^{f},x_{j}^{f}).$ (5)
[…]
For bivariate Normal distributions $x_{i}$ and $x_{j}$, Mutual Information can be measured by $I(x_{i};x_{j})=-\frac{1}{2}\log(1-\rho^{2})$, where $\rho=\frac{\sigma_{x_{i},x_{j}}}{\sigma_{x_{i}}\sigma_{x_{j}}}$ is the correlation coefficient [41]. Accordingly, Eq. (6) yields:
$I(\textbf{x}_{i};\textbf{x}_{j})=-\frac{1}{2}\sum_{f=1}^{F}\log(1-\sigma_{x_{i}^{f},x_{j}^{f}}^{2}).$ (8)
@@
$\text{Cov}(\textbf{g}_{i},\textbf{g}_{j})\approx\sum_{f=1}^{F}\sigma_{x_{i}^{f},x_{j}^{f}}(\frac{\partial\textbf{g}_{i}}{\partial x^{f}_{i}}\bigg{|}_{x^{f}_{i}=\mu^{f}_{i}})(\frac{\partial\textbf{g}_{j}}{\partial x^{f}_{j}}\bigg{|}_{x^{f}_{j}=\mu^{f}_{j}}),$ (9)
where $\sigma_{x_{i}^{f},x_{j}^{f}}$ is the covariance between the $f^{th}$ feature of $x_{i}$ and $x_{j}$. The proof for this Lemma is provided through the Taylor expansion derived in Appendix A.1.
@@
For all clients trained under the federated unsupervised domain generalization setup using self-supervised learning described in Assumption 1, and is extra two clients $C_{i}$ and $C_{j}$, for both positive and negative contrastive data pairs, the sign of $(\frac{\partial\textbf{g}_{i}}{\partial x^{f}_{i}}\bigg{|}_{x^{f}_{i}=\mu^{f}_{i}})(\frac{\partial\textbf{g}_{j}}{\partial x^{f}_{j}}\bigg{|}_{x^{f}_{j}=\mu^{f}_{j}})$ is always positive. See proof in Appendix A.2.
@@
Our theoretical findings may also be applicable to Federated Supervised Domain Generalization, which could form the subject of future research (see Appendix A.5). We further investigate the relationship between the distribution of the domains and their corresponding gradients through the following Corollary.
@@
Given the assumptions stated, for two distinct domains characterized by random variables $\textbf{x}_{i}$ and $\textbf{x}_{j}$ from the distribution family $\mathcal{P}$, as $I(\textbf{x}_{i},\textbf{x}_{j})$ decreases, then the variance of the difference of the corresponding gradients, $\text{Var}(\textbf{g}_{i}-\textbf{g}_{j})$, increases. The full proof is presented in Appendix A.3.

### v2

@@ Abstract
We address the problem of federated domain generalization in an unsupervised setting for the first time. We first theoretically establish a connection between domain shift and alignment of gradients in unsupervised federated learning and show that aligning the gradients at both client and server levels can facilitate the generalization of the model to new (target) domains. Building on this insight, we propose a novel method named FedGaLA, which performs gradient alignment at the client level to encourage clients to learn domain-invariant features, as well as global gradient alignment at the server to obtain a more generalized aggregated model. To empirically evaluate our method, we perform various experiments on four commonly used multi-domain datasets, PACS, OfficeHome, DomainNet, and TerraInc. The results demonstrate the effectiveness of our method which outperforms comparable baselines. Ablation and sensitivity studies demonstrate the impact of different components and parameters in our approach. The source code is available at: https://github.com/MahdiyarMM/FedGaLA.
@@ 1 Introduction
Federated learning (McMahan et al. 2017; Zhang et al. 2021a) has emerged as a promising framework for training machine learning models across multiple decentralized clients while preserving data privacy. It allows clients to collaboratively train a global model without the need to exchange their sensitive and local data. Each client trains a local model using its data and a server aggregates these models at a certain frequency (Ghosh et al. 2020; Charles et al. 2021). However, given that each client collects a different set of local training data, two issues arise. First, the data collected by each client is often recorded under unique conditions that may result in mutual domain shifts (Liu et al. 2021). Second, labeling training data is inherently challenging and resource-intensive; this issue is even more pronounced in the context of federated settings. A typical example of this scenario is a network of wearable activity monitors where variations in user conditions such as demographics or ambient factors can lead to significant domain shifts across devices, meanwhile, the users are generally not asked to provide ground-truth labels for their performed activities.
In prior works, each of these issues has been addressed as a separate problem statement: (i) federated domain generalization (Zhang et al. 2021b; Nguyen, Torr, and Lim 2022; Bai, Bagchi, and Inouye 2024), and (ii) federated unsupervised learning (Zhuang et al. 2021; Han et al. 2022). Despite the effectiveness of both problem definitions, each ignores the fundamental assumptions of the other regarding the data in terms of distributions and availability of labels. To further approach federated learning in a more practical scenario, we propose to merge these two under a new umbrella called federated unsupervised domain generalization, which we define as Definition 1. To our knowledge, prior works have not studied federated learning under such constraints.
@@
To address this new problem, we propose a novel method called Federated Unsupervised Domain Generalization using Global and Local Alignment of Gradients (FedGaLA). To learn more generalized representations from multiple domains, FedGaLA relies on gradient alignment at both client (local) and server (global) levels. At the client level, local models are trained with unlabeled local data available to each client using self-supervised learning (SSL). To learn domain-invariant representations, gradients that are not aligned with the reference gradient, i.e., the global learning direction, are discarded. At the server level, to achieve better generalization, the local models are aggregated based on their alignment with each other (see Figure 1). Specifically, the local gradients that are more closely aligned with the average gradient are given greater weight in the aggregation stage. Moreover, we provide a detailed theoretical framework establishing a connection between the alignment of gradients from different clients and the similarity between their data distributions. Since FedGaLA does not require the sharing of data between clients to perform alignment at either the server or client sides, it effectively preserves data privacy. We verify the performance of our approach using four public datasets, PACS (Yu et al. 2022), OfficeHome (Venkateswara et al. 2017), DomainNet (Peng et al. 2019), and TerraInc (Beery, Van Horn, and Perona 2018) and demonstrate strong performance in federated unsupervised domain generalization in comparison to various baselines. We also conduct different ablation and sensitivity studies to understand the impact of different parameters and the choice of SSL frameworks.
In summary, the main contributions of this study are as follows. (1) To our knowledge, this is the first work to address federated unsupervised domain generalization, introducing a new problem class in this area. (2) We propose a novel technique that aligns gradients at both the local and global levels. In doing so, our solution effectively extracts domain-invariant information in local training and aligns each client’s contributions during the aggregation process, boosting the model’s generalization across various target domains. (3) We conduct extensive experiments and ablations to demonstrate the effectiveness of our proposed method on various benchmarks. To enable fast reproducibility and contribute to the area, we make our code public at: https://github.com/MahdiyarMM/FedGaLA.
@@ 2 Related work
Federated learning. Federating learning is a technique of distributed training that enables learning from decentralized clients without the need to transfer raw data to a central server (McMahan et al. 2017; Zhang et al. 2021a). In the first work in the area of federated learning, FedAVG (McMahan et al. 2017) aggregates the weights of clients via averaging to form the global model. Various federated learning techniques have been since developed for different purposes. FedProx (Li et al. 2020) adds a proximal term to the objective to tackle the data heterogeneity problem. FedBuf (Nguyen et al. 2022) proposes to use buffered asynchronous aggregation to address scalability and privacy issues. FedALA (Zhang et al. 2023b) dynamically aggregates the downloaded global model with the local model on each client, ensuring alignment with the local objective. Moon (Li, He, and Song 2021) leverages the similarity between model representations to improve the local training of individual parties through contrastive learning. FLTrust (Cao et al. 2020) and SignGuard (Xu et al. 2022) leverage gradient filtering to detect and eliminate malicious gradients with the aim of enhancing the robustness of federated learning systems against model poisoning attacks.
@@
Federated domain generalization. This category of problems aims to learn a general model in a privacy-preserving manner that generalizes well to new target domains with distribution shifts (Liu et al. 2021; Li et al. 2023). There have been several studies to explore this direction, mainly focusing on learning domain-invariant features (Nguyen, Torr, and Lim 2022; Wu and Gong 2021) or identifying common features across multiple domains (Zhang et al. 2021b). FedDG (Liu et al. 2021) allows clients to share their data in the frequency space for medical image segmentation. FedSR (Nguyen, Torr, and Lim 2022) proposes to learn a simple representation of data to better generalize to target domains through two regularizers, i.e., L2-norm and conditional Mutual Information. FedADG (Zhang et al. 2021b) leverages adversarial training to align source domain distributions by matching each distribution with a reference distribution. FedKA (Sun, Chong, and Ochiai 2023) learns domain-invariant features by employing feature distribution matching in a universal workspace. CCST (Chen et al. 2023) aligns various client distributions and mitigates model biases by adapting local models to diverse sample styles via cross-client style transfer. FedDG-GA (Zhang et al. 2023c) uses domain divergence and a moment mechanism to enhance generalization through dynamic domain weight adjustment.
FedSB (Soltany et al. 2024) enhances domain generalization by introducing controlled uncertainty to local clients and ensuring balanced contributions from each client. StableFDG (Park et al. 2024) uses style and attention-based strategies to address the federated domain generalization problem. hFedF (Bartholet et al. 2024) utilizes hypernetworks for non-linear aggregation to facilitate generalization to unseen domains. PerAda (Xie et al. 2024) uses knowledge distillation to align each client’s adapter with a global adapter, blending their knowledge. This approach helps the model handle shifting distributions across diverse domains while keeping computational costs low. gPerXAN (Le et al. 2024) combines a personalized normalization scheme with a guiding regularizer to address domain shifts in federated learning, reducing privacy risks and communication costs. MCGDM (Wei and Han 2024) uses intra- and inter-domain gradient matching to minimize domain-specific overfitting, promoting robust generalization on unseen domains. Finally, FGGP (Wan, Huang, and Ye 2024) captures domain information using clustered prototypes and contrastive learning.
@@
Federated unsupervised learning. This category attempts to learn representations from unlabeled data distributed across clients while preserving data privacy (Jin et al. 2020). The study by (van Berlo, Saeed, and Ozcelebi 2020) is the first to propose federated unsupervised learning using an encoder-decoder structure. FedU (Zhuang et al. 2021) employs a contrastive learning approach with online and target networks to enable each client to learn representations from unlabeled data independently. It also introduces a dynamic aggregation mechanism to update the predictor, either locally or globally. Similarly, FedEMA (Zhuang, Wen, and Zhang 2022) uses the exponential moving average of the global model to update local ones. FedX (Han et al. 2022) proposes to use knowledge distillation to learn representation from local data and refine the central server’s knowledge. FedCA (Zhang et al. 2023a) consists of two components: (i) a dictionary module to gather the representations of samples from each client to maintain consistency in the representation space across all clients, and (ii) an alignment module to adjust each client’s representation to match a base model trained on public data.
@@ 3.1 Problem formulation
To formalize Definition 1, assume $K$ clients, $C_{i}$, in a federated setup, each with its own unlabeled data $\mathcal{D}_{i}=\{\textbf{x}_{i}^{(n)}\}_{n=1}^{N_{i}}$. Each dataset consists of $N_{i}$ data points sampled from a distinct data distribution $p(\textbf{x}_{i})$, where $\textbf{x}_{i}$ is a vector of $F$ features, i.e., $\textbf{x}_{i}=[x_{i}^{1},x_{i}^{2},...x_{i}^{F}]^{T}$. The data distributions are assumed to be different among the clients with each distribution $p(\textbf{x}_{i})$ sampled from a family of distributions $\mathcal{P}$. Privacy constraints prevent the transfer of data between clients or to the server $S$. The objective is to learn generalized representations from $\textbf{x}_{i}$ that perform well across unseen distributions $p(\textbf{x}_{t})\sim\mathcal{P}$, where $p(\textbf{x}_{i})\neq p(\textbf{x}_{t})$. This is formulated as minimizing the expected loss over the unseen distributions:
@@
$\min_{\theta_{i}}\mathbb{E}_{p(\textbf{x}_{i})}\left[l_{i}(\theta_{i};\textbf{x})\right]\approx\frac{1}{N_{i}}\sum^{N_{i}}_{n=1}\mathcal{L}(\theta_{i};\textbf{x}_{i}^{(n)}),$ (2)
@@ 3.2 Gradient alignment and domain shift
Under a federated learning framework, privacy constraints prevent clients and servers from accessing each other’s data, including distribution information such as data means and variances. They can, however, observe individual client gradients at the server level and the average aggregated gradient across clients at the client level. We motivate our work on the fact that alignment of gradients may infer characteristics of the client domain distributions, thus facilitating improved model generalization. While empirically the utility of gradients has been demonstrated in the area of domain generalization (Mansilla et al. 2021), no theoretical basis has been proposed for this approach under federated constraints. Accordingly, we aim to establish a link between gradient alignment and domain shifts within the proposed problem formulation (Definition 1). Our theoretical findings provide the basis for effective local parameter updates and global model aggregation to address federated unsupervised domain generalization.
@@
Let each $\textbf{x}_{i}^{f}$ be a random variable drawn from a Normal distribution $p(\textbf{x}_{i}^{f})\sim\mathcal{P}$ (Bar-Hillel et al. 2003). Within a single domain, following (Štrumbelj and Kononenko 2014, 2011), features are assumed to be independent ($Cov(x_{i}^{f_{1}},x_{i}^{f_{2}})=0$). Across different domains, corresponding features of $\textbf{x}_{i}^{f}$ and $\textbf{x}_{j}^{f}$ are bivariate with a covariance of $\sigma_{x_{i}^{f},x_{j}^{f}}$. In line with contemporary practices in deep learning, the features are normalized with $\mu=0$ and $\sigma^{2}=1$ (Yu and Spiliopoulos 2022; Qi, Zhou, and Wang 2022; Huang et al. 2023). We also assume that each client is a one-layer encoder with the sigmoid activation function. Following (Chen et al. 2020), local training is performed using contrastive loss where a random augmentation of a data sample is used as the positive pair and all other samples are utilized as the negative pairs. The random augmentation is performed with the same random Affine transformation ($Ax_{i}+B$) (Wang et al. 2023) broadcast to all clients. After each epoch, the local models are aggregated in the server and sent back to all clients. The gradient $\textbf{g}_{i}$ of the model is assumed to be differentiable.
@@
Given Assumption 1, under the problem proposed in Definition 1, for two distinct domains characterized by random variables $\textbf{x}_{i}$ and $\textbf{x}_{j}$ belonging to two different clients $C_{i}$ and $C_{j}$, an increase in domain shift across the clients results in a decrease in covariance $\text{Cov}(\emph{$\textbf{g}_{i}$},\emph{$\textbf{g}_{j}$})$ of the corresponding gradients $\textbf{g}_{i}$, $\textbf{g}_{j}$ across $C_{i}$ and $C_{j}$’s respective local models.
@@
Motivated by previous works where the similarity between the representations of different domains under domain shift is measured by Mutual Information (Gao et al. 2020; Menapace, Lathuilière, and Ricci 2020; Wang et al. 2021), we use this concept for modeling the similarity between different domains drawn from a family of distributions. To calculate the mutual information we introduce the following Lemma.
Lemma 1.
Given Assumption 1, the mutual information between two random variables $\textbf{x}_{i}$ and $\textbf{x}_{j}$ can be calculated as:
$I(\textbf{\emph{x}}_{i};\textbf{\emph{x}}_{j})=-\frac{1}{2}\sum_{f=1}^{F}\log(1-\sigma_{x_{i}^{f},x_{j}^{f}}^{2}).$ (3)
The full proof of this lemma is presented in Appendix A.1.
@@
$\emph{\text{Cov}}(\emph{{g}}_{i},\textbf{\emph{g}}_{j})_{mn}\approx\sum_{f=1}^{F}\sigma_{x^{f}_{i},x^{f}_{j}}\left(\frac{\partial g_{i(m)}}{\partial x_{i}^{f}}\bigg{|}_{\mu_{i}^{f}}\right)\left(\frac{\partial g_{j(n)}}{\partial x_{j}^{f}}\bigg{|}_{\mu_{j}^{f}}\right),$ (4)
where $\sigma_{x_{i}^{f},x_{j}^{f}}$ is the covariance between the $f^{th}$ feature of $x_{i}$ and $x_{j}$ and ${g}_{i(m)}$ is the $m^{th}$ dimension of $g_{i}$. The proof for this Lemma is provided through the Taylor expansion derived in Appendix A.2.
@@
For all clients trained under the federated unsupervised domain generalization setup using self-supervised learning described in Assumption 1, for two clients $C_{i}$ and $C_{j}$, for both positive and negative contrastive data pairs, the sign of $(\frac{\partial\textbf{\emph{g}}_{i}}{\partial x^{f}_{i}}\bigg{|}_{x^{f}_{i}=\mu^{f}_{i}})(\frac{\partial\textbf{\emph{g}}_{j}}{\partial x^{f}_{j}}\bigg{|}_{x^{f}_{j}=\mu^{f}_{j}})$ is always positive in all dimensions. See proof in Appendix A.3.
@@
Our theoretical findings may also be applicable to Federated Supervised Domain Generalization, which could form the subject of future research as expressed in Appendix A.4 and (Molahasani et al. 2024). We further investigate the relationship between the distribution of the domains and their corresponding gradients through the following Corollary.
@@
Given the assumptions stated, for two distinct domains characterized by random variables $\textbf{\emph{x}}_{i}$ and $\textbf{\emph{x}}_{j}$ from the distribution family $\mathcal{P}$, as $I(\textbf{\emph{x}}_{i},\textbf{\emph{x}}_{j})$ decreases, then the variance of the difference of the corresponding gradients, $\emph{\text{Var}}(\textbf{\emph{g}}_{i}-\textbf{\emph{g}}_{j})$, increases. The full proof is presented in Appendix A.6.